# Almost sure convergence rates of stochastic gradient methods under gradient domination

### v1

@@ Appendix F Proof of Section 5
$\tilde{D}_{n+1}$ $\leq(1-\gamma_{n}c^{2})\tilde{D}_{n}+\gamma_{n}\xi_{n}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2},$ (43)
$\leq\tilde{D}_{1}\prod_{k=1}^{n}(1-\gamma_{k}c^{2})+\sum_{k=1}^{n}\gamma_{k}\xi_{k}\mathbf{1}_{\Omega_{k}}+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2}\mathbf{1}_{\Omega_{k}}.$ (44)
@@
$\tilde{D}_{n+1}$ $\leq(1-\gamma_{n}^{q}c^{2})\tilde{D}_{n}+(2\beta)^{-\frac{1}{2\beta-1}}(1-\frac{1}{2\beta}){c^{2}}\gamma_{n}^{\frac{2\beta q-1}{2\beta-1}}+\gamma_{n}\xi_{n}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\lVert V_{n+1}(X_{n})\rVert^{2}$ (45)
$\leq\tilde{D}_{1}\prod_{k=1}^{n}(1-\gamma_{k}^{q}c^{2})+\tilde{c}\sum_{k=1}^{n}\gamma_{k}^{\frac{2\beta q-1}{2\beta-1}}+\sum_{k=1}^{n}\gamma_{k}\xi_{k}\mathbf{1}_{\Omega_{k}}+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2},$ (46)
@@
The proof follows line for line as in Lemma F.4 by replacing $l$ with $f^{\ast}$ and taking the different definition of $\tilde{D}_{n}$ and $\Omega_{n}$ into account. ∎
@@
$M_{n}=\sum_{k=1}^{n}\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{k}}\quad\textrm{ and }\quad S_{n}=\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2}\mathbf{1}_{\Omega_{k}}.$
Then, $(M_{n})_{n\in\mathbb{N}}$ is a $(\mathcal{F}_{n-1})$-martingale with zero mean and $(S_{n})_{n\in\mathbb{N}}$ is a $(\mathcal{F}_{n-1})$-sub-martingale by Assumption 2.4. Note that by the choice of $\gamma_{n}$ we have that $\sum_{n}\gamma_{n}^{2}<\infty$ and hence $\mathbb{E}[S_{n}]<\infty$ for all $n\in\mathbb{N}$. Next, define $R_{n}=M_{n}^{2}+S_{n}$ for every $n\in\mathbb{N}$. Moreover, let
@@
which is an $\mathcal{F}_{n}$-measurable event on $(\Omega,\mathcal{F},\mathbb{P})$. We define $R_{0}=0$ such that $E_{0}=\Omega$.
With these definitions we can directly prove a parallel result to Lemma F.6 without the auxiliary result in Lemma F.5.
Lemma F.10.
@@
b) Note that $E_{0}=\Omega=\Omega_{1}$ because $X_{1}\in\mathcal{U}_{1}=\Omega_{1}$ almost surely. We prove the assertion by induction. Let $\omega\in E_{n}$. Since $E_{n}\subset E_{n-1}\subset\Omega_{n}$ by induction assumption, we have $\omega\in\Omega_{n}$ and thus $\omega\in\Omega_{k}$ for all $k\leq n$. It remains to show that $X_{n+1}(\omega)\in\mathcal{U}$ to prove that $\omega\in\Omega_{n+1}$. We separate both cases for $\beta$:
$\beta=\frac{1}{2}$: The inequality (43) and the induction hypothesis yield
$D_{n+1}(\omega)$ $\leq D_{1}(\omega)\prod_{k=1}^{n}(1-\gamma_{k}c)+\sum_{k=1}^{n}\gamma_{k}\xi_{k+1}(\omega)\mathbf{1}_{\Omega_{k}}(\omega)+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k}(\omega))\rVert^{2}$
@@
$\beta\in(\frac{1}{2},1]$: Similarly, we obtain from (45)
$D_{n+1}(\omega)$ $\leq D_{1}(\omega)\prod_{k=1}^{n}(1-\gamma_{k}^{q}c)+\tilde{c}\sum_{k=1}^{n}\gamma_{k}^{\frac{2\beta q-1}{2\beta-1}}+\sum_{k=1}^{n}\gamma_{k}\xi_{k+1}(\omega)\mathbf{1}_{\Omega_{k}}(\omega)+\frac{L}{2}\sum_{k=0}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k}(\omega))\rVert^{2}$
@@
c) Follows line by line as in Lemma F.6 part c).
Lemma F.11.
Let $\delta>0$ be a tolerance level and $\gamma_{n}\leq\gamma_{1}$ be sufficiently small such that $\sum_{n=1}^{\infty}\gamma_{n}^{2}<\frac{\delta\epsilon}{2(G^{2}C^{2}+G^{2}+C)}$ and the condition in Lemma F.10 is fulfilled. Then, we have
@@
Line by line as in Lemma F.7. ∎
@@
Hence, using Lemma F.11 we obtain
@@
Appendix G Proofs of Section 7
@@
Case $\lambda=0$:
Define the optimal reward gap in every state $s\in\mathcal{S}$ by
@@
where $a^{\ast}(s)$ denotes the best possible action in state $s$ and $Q^{\ast}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ denotes the optimal Q-function defined by $Q^{\ast}(s,a)=\mathbb{E}_{\mu}^{\pi^{\ast}}[\sum_{t=0}^{\infty}\rho^{-t}r(S_{t},A_{t})|A_{0}=a]$. W.l.o.g. we assume that $a^{\ast}(s)$ is unique. Similary let $Q^{\pi_{w}}(s,a)=\mathbb{E}_{\mu}^{\pi_{w}}[\sum_{t=0}^{\infty}\rho^{-t}r(S_{t},A_{t})|A_{0}=a]$ be the Q-function for policy $\pi_{w}$.
@@
Case $\lambda>0$:
For any $\alpha\in(0,1)$ choose $r=\alpha^{2}\exp\left(\frac{-1}{(1-\rho)\lambda}\right)^{2}\frac{\lambda\min_{s}\mu(s)}{2\mathrm{ln}2}$ and assume that $w\in\mathcal{B}_{r,\lambda}^{\ast}$. By [10, Lem. 12] we have
@@
where the last inequality is due to [37, Thm. 1]. It follows directly that

### v3

@@ Appendix D Proofs of Section 5
$\begin{split}{D}_{n+1}\mathbf{1}_{\Omega_{n}}&\leq(1-\gamma_{n}c^{2})D_{n}\mathbf{1}_{\Omega_{n}}+\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2},\\ &\leq{D}_{1}\prod_{k=1}^{n}(1-\gamma_{k}c^{2})\mathbf{1}_{\Omega_{n}}+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}c^{2})\right)\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{n}}\\ &\quad+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2}\mathbf{1}_{\Omega_{n}}.\end{split}$ (41)
@@
$\begin{split}{D}_{n+1}\mathbf{1}_{\Omega_{n}}&\leq(1-\gamma_{n}^{q}c^{2}){D}_{n}\mathbf{1}_{\Omega_{n}}+(2\beta)^{-\frac{1}{2\beta-1}}(1-\frac{1}{2\beta}){c^{2}}\gamma_{n}^{\frac{2\beta q-1}{2\beta-1}}+\gamma_{n}\xi_{n}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\lVert V_{n+1}(X_{n})\rVert^{2}\\ &\leq{D}_{1}\prod_{k=1}^{n}(1-\gamma_{k}^{q}c^{2})\mathbf{1}_{\Omega_{n}}+\tilde{c}\sum_{k=1}^{n}\gamma_{k}^{\frac{2\beta q-1}{2\beta-1}}+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}^{q}c^{2})\right)\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{n}}\\ &\quad+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2}\mathbf{1}_{\Omega_{n}},\end{split}$ (42)
@@
The proof follows line for line as in Lemma D.4 by replacing $l$ with $f^{\ast}$ and taking the different definition of $\tilde{D}_{n}$ and $\Omega_{n}$ into account. ∎
@@
$M_{n}=\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}c^{2})\right)\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{k}},\quad M_{n}^{(q)}=\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}^{q}c^{2})\right)\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{k}}\quad$
[…]
Then, $(M_{n})_{n\in\mathbb{N}}$ and $(M_{n}^{(q)})$ are $(\mathcal{F}_{n+1})$-martingales with zero mean and $(S_{n})_{n\in\mathbb{N}}$ is a $(\mathcal{F}_{n+1})$-sub-martingale by Assumption 2.4. Note that by the choice of $\gamma_{n}$ we have $\sum_{n}\gamma_{n}^{2}<\infty$ and hence $\mathbb{E}[S_{n}]<\infty$ for all $n\in\mathbb{N}$. Next, define $R_{n}=M_{n}^{2}+S_{n}$ and $R_{n}=(M_{n}^{(q)})^{2}+S_{n}$ respectively (with some abuse of notation) for every $n\in\mathbb{N}$. Moreover, let
@@
which is a $\mathcal{F}_{n+1}$-measurable event on $(\Omega,\mathcal{F},\mathbb{P})$. We define $R_{0}=0$ such that $E_{0}=\Omega$.
With these definitions, we can directly prove a parallel result to Lemma D.6 without the auxiliary result in Lemma D.5.
Lemma D.10.
@@
b) Note that $E_{0}=\Omega=\Omega_{1}$ because $X_{1}\in\mathcal{U}_{1}=\Omega_{1}$ almost surely. We prove the claim by induction. Let $\omega\in E_{n}$. Since $E_{n}\subset E_{n-1}\subset\Omega_{n}$ by induction assumption, we have $\omega\in\Omega_{n}$ and thus $\omega\in\Omega_{k}$ for all $k\leq n$. It remains to show that $X_{n+1}(\omega)\in\mathcal{U}$ to prove that $\omega\in\Omega_{n+1}$. We separate both cases for $\beta$:
$\beta=\frac{1}{2}$: The inequality Equation 41 and the induction hypothesis yield
$D_{n+1}(\omega)$ $\leq D_{1}(\omega)\prod_{k=1}^{n}(1-\gamma_{k}c)+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}c^{2})\right)\gamma_{k}\xi_{k+1}(\omega)+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k}(\omega))\rVert^{2}$
$=D_{1}(\omega)\prod_{k=1}^{n}(1-\gamma_{k}c)+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}c^{2})\right)\gamma_{k}\xi_{k+1}(\omega)\mathbf{1}_{\Omega_{k}}(\omega)$
$\quad+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k}(\omega))\rVert^{2}\mathbf{1}_{\Omega_{k}}(\omega)$
@@
$\beta\in(\frac{1}{2},1]$: Similarly, we obtain from Equation 42
$D_{n+1}(\omega)$ $\leq D_{1}(\omega)\prod_{k=1}^{n}(1-\gamma_{k}^{q}c)+\tilde{c}\sum_{k=1}^{n}\gamma_{k}^{\frac{2\beta q-1}{2\beta-1}}+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}^{q}c^{2})\right)\gamma_{k}\xi_{k+1}(\omega)$
$\quad+\frac{L}{2}\sum_{k=0}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k}(\omega))\rVert^{2}$
$=D_{1}(\omega)\prod_{k=1}^{n}(1-\gamma_{k}^{q}c)+\tilde{c}\sum_{k=1}^{n}\gamma_{k}^{\frac{2\beta q-1}{2\beta-1}}+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}^{q}c^{2})\right)\gamma_{k}\xi_{k+1}(\omega)\mathbf{1}_{\Omega_{k}}(\omega)$
$\quad+\frac{L}{2}\sum_{k=0}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k}(\omega))\rVert^{2}\mathbf{1}_{\Omega_{k}}(\omega)$
@@
c) Follows line by line as in Lemma D.6 part c). ∎
Lemma D.11.
Let $\delta>0$ be a tolerance level and $\gamma_{n}\leq\gamma_{1}$ be sufficiently small such that $\sum_{n=1}^{\infty}\gamma_{n}^{2}<\frac{\delta\epsilon}{2(G^{2}C^{2}+G^{2}+C)}$ and the condition in Lemma D.10 is fulfilled. Then, we have
@@
Line by line as in Lemma D.7. ∎
@@
Hence, using Lemma D.11 we obtain
@@
Appendix E Proofs of Section 7
In the following section, we provide the proof of Section 7.
@@ Appendix E Proofs of Section 7
Case $\lambda=0$: Define the optimal reward gap in every state $s\in\mathcal{S}$ by
@@
where $a^{\ast}(s)$ denotes the best possible action in state $s$ and $Q^{\ast}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ denotes the optimal Q-function defined by $Q^{\ast}(s,a)=\mathbb{E}_{\mu}^{\pi^{\ast}}[\sum_{t=0}^{\infty}\rho^{-t}r(S_{t},A_{t})|A_{0}=a]$. W.l.o.g. we assume that $a^{\ast}(s)$ is unique. Similarly let $Q^{\pi_{w}}(s,a)=\mathbb{E}_{\mu}^{\pi_{w}}[\sum_{t=0}^{\infty}\rho^{-t}r(S_{t},A_{t})|A_{0}=a]$ be the Q-function for policy $\pi_{w}$.
@@
Case $\lambda>0$: For any $\alpha\in(0,1)$ choose $r=\alpha^{2}\exp\left(\frac{-1}{(1-\rho)\lambda}\right)^{2}\frac{\lambda\min_{s}\mu(s)}{2\mathrm{ln}2}$ and assume that $w\in\mathcal{B}_{r,\lambda}^{\ast}$. By Ding et al. [10, Lem. 12] we have
@@
where the last inequality is due to Nachum et al. [40, Thm. 1]. It follows directly that